OpenStudy (anonymous):

If a tightrope walker falls, he will land on a safety net. His height h in feet after a fall can be modeled by h(t) = 60 - 16t2, where t is the time in seconds. If the safety net is 11 feet above the ground, how many seconds will the tightrope walker fall before landing on the safety net? Round your answer to the nearest hundredth. _____ s

OpenStudy (lgbasallote):

hint: 60 - 16t^2 = 11 solve for t does that help?

OpenStudy (anonymous):

Do you subtract like 60 or 11 from both sides?

OpenStudy (lgbasallote):

yes. you subtract 60 from both sides

OpenStudy (anonymous):

-16^2=-49 right? and then would you square both sides?

OpenStudy (anonymous):

-16t^2=-49

OpenStudy (lgbasallote):

no...you divide both sides by -16

OpenStudy (anonymous):

so t^2=3.0625?

OpenStudy (lgbasallote):

don't use a calculator

OpenStudy (lgbasallote):

put them in fraction form

OpenStudy (anonymous):

-16/16t^2=-49/16?

OpenStudy (lgbasallote):

divide by -16...not just 16

OpenStudy (anonymous):

the first part would be 1 though right?

OpenStudy (lgbasallote):

right. what about the right side?

OpenStudy (anonymous):

3 1/16?

OpenStudy (lgbasallote):

actually it's 49/16 \[\implies t^2 = \frac{49}{16}\] make sense?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

now take the square root of both sides

OpenStudy (anonymous):

t= 7/4?

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now put that in decimal form

OpenStudy (anonymous):

t=1.75

OpenStudy (lgbasallote):

right.

OpenStudy (lgbasallote):

so your answer is 1.75 seconds

OpenStudy (anonymous):

thank you, for your help

OpenStudy (lgbasallote):

welcome